OpenStudy (anonymous):

Lisa surveyed 60 students at her school and found that 0.85 of the students she surveyed said their favorite class is math. Another 15% of the students she surveyed reported that their favorite class is science. How many more students in the survey prefer math over science? 9 15 35 42

OpenStudy (anonymous):

@iambatman @Michele_Laino

OpenStudy (michele_laino):

please, note that to each percentage correspond a precise number of students

OpenStudy (anonymous):

15% = 15 people

OpenStudy (michele_laino):

for example 15% * 20 people = 3 people

OpenStudy (anonymous):

0.85 = 85

OpenStudy (anonymous):

so then...

OpenStudy (michele_laino):

you have to multiply your percentage by the total number of students

OpenStudy (anonymous):

15 * 60 = 900

OpenStudy (michele_laino):

not 15, but 15/100=0.15

OpenStudy (anonymous):

so then .85 - .15 = .70?

OpenStudy (michele_laino):

no, please no 15*60, but 0.15*60=...

OpenStudy (anonymous):

= 9

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

so then its a?

OpenStudy (michele_laino):

now, do the same using other percentage, please!

OpenStudy (anonymous):

= 51

OpenStudy (michele_laino):

well done!

OpenStudy (anonymous):

42?

OpenStudy (anonymous):

so d

OpenStudy (michele_laino):

perfect!

OpenStudy (anonymous):

If a sprinkler waters 1 over 20 of a lawn in 1 over 5 of an hour, how much time will it take to water the entire lawn? 3 over 20 hour 1 over 4 hour 4 hours 5 hours

OpenStudy (anonymous):

d?

OpenStudy (michele_laino):

every 1 over 5 of hour our sprinkler waters 1 over 20 of lawn, so after 2*(1 over 5) hour our sprinkler waters 2*( 1 over 20) of lawn and after 3*(1 over 5) hour, our sprinkler waters 3*(1 over 20) lawn, and so on... so?

OpenStudy (michele_laino):

20*(1 over 20) lawn are watered after 20 *(1 over 5) hour, so?

OpenStudy (michele_laino):

please try!

OpenStudy (anonymous):

so would it be b? i did try!!

OpenStudy (michele_laino):

no, please what is 20*(1 over 5)?

OpenStudy (anonymous):

4

OpenStudy (michele_laino):

that's right!

OpenStudy (michele_laino):

so?

OpenStudy (anonymous):

oh so then its c... thanks!!!

OpenStudy (michele_laino):

ok! thanks!!